# Optimal Attention Temperature Enhances In-Context Learning under Distribution Shift

### v1

@@ Abstract
Pretrained Transformers excel at in-context learning (ICL), inferring new tasks from only a handful of examples. Yet, their ICL performance can degrade sharply under distribution shift between pretraining and test data—a regime increasingly common in real-world deployments. While recent empirical work hints that adjusting the attention temperature in the softmax can enhance Transformer performance, the attention temperature’s role in ICL under distribution shift remains unexplored. This paper provides the first theoretical and empirical study of attention temperature for ICL under distribution shift. Using a simplified but expressive “linearized softmax” framework, we derive closed-form generalization error expressions and prove that shifts in input covariance or label noise substantially impair ICL, but that an optimal attention temperature exists which minimizes this error. We then validate our predictions through extensive simulations on linear regression tasks and large-scale experiments with GPT-2 and LLaMA2-7B on question-answering benchmarks. Our results establish attention temperature as a principled and powerful mechanism for improving the robustness of ICL in pretrained Transformers—advancing theoretical understanding and providing actionable guidance for selecting attention temperature in practice.
@@ 1 Introduction
where $\bm{Z}$} denotes the input representation and $\bm{Q}$}, $\bm{K}$}, and $\bm{V}$} are the query, key, and value weight matrices, respectively. The parameter $\tau>0$, referred to as the attention temperature, controls the variance of the softmax outputs and hence the selectivity of attention weights. This quantity is distinct from the “sampling temperature” commonly used to adjust the output distribution of generative models such as large language models (LLMs) (Renze & Guven, 2024). Throughout this work, we exclusively focus on the attention temperature as an intrinsic component of the attention mechanism.
@@
In this paper, we present a unified theoretical and empirical study of the attention temperature in the context of in-context learning (ICL). We focus on how adjusting this parameter can systematically improve the ICL performance of pretrained Transformers under distributional shift. We address this question in the setting of linear regression tasks, which offer a well-controlled yet expressive framework for dissecting the mechanisms of ICL (Garg et al., 2022; Zhang et al., 2024). Departing from prior work restricted to linear attention, we analyze a Transformer with linearized softmax attention—an architecture that preserves the essential temperature-dependent behavior of standard attention while remaining mathematically tractable.
@@
    We provide, to our knowledge, the first theoretical characterization of the optimal attention temperature for pretrained Transformers with linearized softmax attention in ICL tasks.
@@ 2 Related work
Simplified Transformer variants—particularly those using linear attention—have proven useful for gaining analytical insights about ICL (Garg et al., 2022; Zhang et al., 2024; Raventós et al., 2023). Notably, Zhang et al. (2024) showed that linear Transformers approximate Bayes-optimal inference in linear regression tasks, even under distribution shift. We build on this line of research but focus explicitly on the role of the attention temperature. In contrast to Zhang et al. (2024), we (i) employ linearized softmax attention to isolate the effect of temperature, (ii) study how temperature adjustments can mitigate the impact of distribution shifts, and (iii) derive and empirically evaluate the optimal temperature for improving ICL performance. These advances extend prior analyses and yield a deeper theoretical and empirical understanding of how principled temperature selection enhances the robustness of Transformers under distributional shift. [^1] Additional related work is discussed in Appendix L.
@@
Although linear attention has gained traction for its computational efficiency, it typically lags behind softmax-based counterparts in predictive performance, spurring efforts to narrow this gap (Choromanski et al., 2021; Qin et al., 2022). A pivotal advance in this direction is due to Han et al. (2024), who showed that a linearized variant of softmax attention can closely approximate the performance of standard softmax attention. Building on this insight, we adopt the linearized softmax formulation, which preserves the essential temperature-dependent behavior of standard attention while enabling tractable theoretical analysis. This choice provides a principled framework for investigating how attention-temperature selection shapes ICL performance in pretrained Transformers.
@@ 3 Setting
We describe the setup for analyzing ICL in linear regression using pretrained Transformers, covering the data model, linearized attention with reparameterization, evaluation metrics, and the Bayes-optimal benchmark.
Notation —
3.1 Problem setup: In-context learning for linear regression
@@
${\bm{x}}_{i}\sim{\mathcal{N}}({\bm{\mu}}_{x},{\bm{\Sigma}}_{x}),\quad y_{i}={\bm{w}}^{T}{\bm{x}}_{i}+\epsilon_{i},\quad\epsilon_{i}\sim{\mathcal{N}}(0,\sigma^{2}),$ (2)
@@
$\|{\bm{\mu}}_{x}\|,\|{\bm{\mu}}_{w}\|\leq c_{1},\quad\lambda_{\min}({\bm{\Sigma}}_{x}),\lambda_{\min}({\bm{\Sigma}}_{w})\geq c_{2},\quad\lambda_{\max}({\bm{\Sigma}}_{x}),\lambda_{\max}({\bm{\Sigma}}_{w})\leq c_{3}.$
@@ 3.2 Modeling attention with transformers
3.3 Linearized attention approximation
To analytically characterize the effect of temperature on ICL, we employ a linearized approximation of softmax attention (see Appendix B for the derivation and formal definition):
${\bm{E}}:={\bm{Z}}+\frac{1}{l}{\bm{V}}{\bm{Z}}\left(\frac{({\bm{K}}{\bm{Z}})^{T}({\bm{Q}}{\bm{Z}})}{\tau}+\mathbf{1}-\frac{\mathbf{1}}{l}\sum_{j=1}^{l}\frac{({\bm{K}}{\bm{Z}}_{:,j})^{T}({\bm{Q}}{\bm{Z}})}{\tau}\right),$ (5)
where $\hat{y}:=E_{d+1,l}$ represents the predicted label. In contrast to linear attention (Zhang et al., 2024),
@@ 3.3 Linearized attention approximation
Remark 3.4 (Linear vs. linearized attention).
Linearized attention maintains row-wise normalization, making it inherently more robust to shifts in input means — a critical failure mode of linear attention in ICL. Appendix C provides an illustrative comparison.
Another key distinction between linear case and linearized softmax case is that linear (with temperature scaling) fails to capture the temperature behavior of softmax. However, while this may not be immediately apparent, linearized softmax closely mirrors the behavior of softmax with respect to temperature variation. A detailed explanation together with an illustrative example is provided in Appendix D.
3.4 Reparametrization of linearized attention
@@
By analyzing this reparameterization, we gain a deeper understanding of how the model parameters interact with the data to address the ICL problem effectively. This foundational insight will provide the necessary basis for discussing the pretraining of these parameters in Section 4.1.
@@ 3.6 Bayes-optimal ridge estimator
The derivation, provided in Appendix A, illustrates how Bayesian principles inform regression by integrating data evidence with prior distributions to yield more reliable predictions. In our setting, the inputs and labels are derived from the prompt matrix $\bm{Z}$}, and the Bayes-optimal linear model predicts any input $\bm{x}$} as $\hat{{\bm{w}}}_{Bayes}^{T}{\bm{x}}$.
4 Theoretical results
In this section, we present our main theoretical results on the ICL under distribution shifts for the Transformer with a linearized attention without MLP layers, denoted by (5). We begin by showing how to pretrain the model to approximate the Bayes-optimal linear predictor, thereby grounding its predictive performance. We then identify specific conditions under which the model fails to generalize under distribution shifts at test time, revealing key limitations of the model in ICL. Following this, we provide a detailed characterization of its generalization error, offering a principled framework for analyzing performance. Finally, we investigate the role of the temperature parameter and demonstrate that adjusting it appropriately can substantially improve generalization—especially in cases where the model initially fails to perform effective in-context learning.
4.1 Model pretraining
We begin our pretraining analysis by observing that the prediction generated by the Transformer (5) can be reduced to the following form (see Appendix E for the derivation):
$\hat{y}({\bm{Z}};{\bm{V}},{\bm{M}}):=E_{d+1,l}=\frac{1}{\tau}\hat{{\bm{w}}}_{Att}({\bm{C}}_{xx},{\bm{C}}_{xy},C_{yy};{\bm{M}},{\bm{V}})^{T}{\bm{x}}_{l}+b_{Att}({\bm{s}}_{x},s_{y};{\bm{V}}),$ (10)
where $\hat{{\bm{w}}}_{Att}({\bm{C}}_{xx},{\bm{C}}_{xy},C_{yy};{\bm{M}},{\bm{V}})\in\mathbb{R}^{d}$ and $b_{Att}({\bm{s}}_{x},s_{y};{\bm{V}})\in\mathbb{R}$. ${\bm{s}}_{x}$ and $s_{y}$ denote the sample means of the input $\bm{x}$} and the label $y$, respectively, and ${\bm{C}}_{xx}$ and ${\bm{C}}_{xy}$ are the sample covariances corresponding to $\text{Cov}({\bm{x}})$ and $\text{Cov}({\bm{x}},y)$. These statistics are computed from the prompt matrix $\bm{Z}$}.
For pretraining, we optimize the parameters $\bm{V}$} and $\bm{M}$} using $m$ samples of $({\bm{Z}},y_{l})$ drawn from the distribution $\mathcal{D}^{train}$, where each $\bm{Z}$} contains l1l-1 $({\bm{x}},y)$ pairs intended for ICL. Building upon prior work that connects ICL in linear regression to the Bayes-optimal ridge estimator (Zhang et al., 2024; Raventós et al., 2023), we configure $\bm{M}$} and $\bm{V}$} to emulate Bayes-optimal ridge regression. Specifically, we aim for $\hat{{\bm{w}}}_{Att}({\bm{C}}_{xx},{\bm{C}}_{xy};{\bm{M}},{\bm{V}})\approx\hat{{\bm{w}}}_{Bayes}$ and $b_{Att}({\bm{s}}_{x},s_{y};{\bm{V}})\approx 0$.
Lemma 4.1 (Pretrained Parameters).
This lemma establishes a theoretical connection between the pretrained parameters and the Bayes-optimal estimator, reinforcing the foundation of our approach.
Moreover, specific instances of Lemma 4.1 recover settings explored in prior studies. For example, under the assumptions ${\bm{\Sigma}}_{x}={\bm{\Sigma}}_{w}={\bm{I}}$, ${\bm{\mu}}_{w}=\mathbf{0}$, and $\sigma=0$, Von Oswald et al. (2023) employ ${\bm{M}}_{11}=\text{Cov}(x)^{-1}$ and ${\bm{v}}_{21}=\mathbf{0}$ within a linear attention framework. Our formulation generalizes this by allowing ${\bm{v}}_{21}\neq\mathbf{0}$, which reflects our assumption that ${\bm{\mu}}_{w}\neq\mathbf{0}$—a departure from earlier works. Indeed, our analysis reveals that ${\bm{v}}_{21}$ encodes information related to task vector bias ${\bm{\mu}}_{w}$. Additionally, our choice of ${\bm{M}}_{11}$ explicitly accounts for label noise ($\sigma^{2}$), thereby enhancing the model’s adaptability and maintaining a Bayesian interpretation.
We conclude this section with two remarks on task diversity and parameter optimality:
Remark 4.2.
A high degree of task diversity (i.e., the number of distinct tasks) is essential for enabling effective in-context learning (Wu et al., 2024). Within our framework, task diversity directly impacts the accuracy of estimating ${\bm{\mu}}_{w}$ and ${\bm{\Sigma}}_{w}$ during pretraining.
Remark 4.3.
While the pretrained parameters specified in Lemma 4.1 are not guaranteed to be optimal in all settings, they are analytically useful for examining the effects of distribution shifts and the role of the temperature parameter in ICL. Importantly, our characterization of ICL performance and temperature optimality does not depend on these particular parameter choices.
Based on Lemma 4.1, we arrive at the following corollary:
Corollary 4.4.
Suppose there is no distribution shift between training and inference. Then, under the parameter configuration of Lemma 4.1, the Transformer model (5) emulates the Bayes-optimal linear model, implying that it is capable of in-context learning according to Definition $3.$
4.2 Effect of distribution shift
Case I: Shift in input distribution —
  1. (i)
    Mean shift (${\bm{\mu}}_{x}^{train}\neq{\bm{\mu}}_{x}^{test}$): Centering renders the linearized model invariant to mean shifts, but the uncentered linear attention model remains sensitive, as noted in Remark 3.4.
  2. (ii)
Case II: Shift in Task Distribution —
The task vectors follow ${\bm{w}}\sim\mathcal{N}({\bm{\mu}}_{w},{\bm{\Sigma}}_{w})$. Let ${\bm{\mu}}_{w}^{train},{\bm{\Sigma}}_{w}^{train}$ and ${\bm{\mu}}_{w}^{test},{\bm{\Sigma}}_{w}^{test}$ be the mean and covariance of the task distribution during pretraining and testing, respectively. The Transformer model can incorporate ${\bm{\mu}}_{w}^{train}$ and ${\bm{\Sigma}}_{w}^{train}$ via the pretrained parameters ${\bm{M}}_{11}$ and ${\bm{v}}_{21}$ (see Lemma 4.1). However, as the context length $l$ increases, the model’s dependence on the task distribution diminishes. Thus, shifts in the task distribution primarily affect ICL performance for small $l$.
Case III: Shift in noise distribution —
Summary — The Transformer model is robust to shifts in input mean but sensitive to input covariance changes. Shifts in task or noise distribution reduce ICL performance at small $l$, though increasing $l$ mitigates these effects. In Section 4.4, we explore optimal temperature selection as a way to enhance robustness. Before that, we analyze the generalization error of the model in the next section.
@@ 4.3 In-context learning performance
We analyze the in-context learning (ICL) performance of the Transformer model (5) by evaluating the generalization error defined in (8). To establish a general setting for the subsequent results, we impose the following assumption on the pretrained parameters:
@@
Note that the pretrained parameters obtained in Lemma 4.1 satisfy Assumption 4.5 with high probability under Assumptions 3.13.2. However, the generalization error result stated below holds for any parameters ${\bm{M}},{\bm{V}}$ that satisfy Assumption 4.5.
@@
Suppose Assumptions 3.13.2 and 4.5 hold. At test time, assume the input, task, and noise distributions are given by $\mathcal{N}({\bm{\mu}}_{x},{\bm{\Sigma}}_{x})$, $\mathcal{N}({\bm{\mu}}_{w},{\bm{\Sigma}}_{w})$, and $\mathcal{N}(0,\sigma^{2})$, respectively. Define
@@
$\mathcal{G}({\bm{V}},{\bm{M}})$ $=\frac{1}{\tau^{2}}\operatorname{Tr}\left({\bm{A}}{\bm{M}}_{11}^{T}{\bm{F}}_{1}{\bm{M}}_{11}\right)-\frac{1}{\tau}\operatorname{Tr}\left({\bm{A}}\left({\bm{F}}_{2}{\bm{M}}_{11}+{\bm{M}}_{11}^{T}{\bm{F}}_{2}^{T}\right)\right)+\operatorname{Tr}\left({\bm{A}}{\bm{B}}\right)+\sigma^{2},$ (12)
@@
${\bm{F}}_{1}$ $:=\left({\bm{\Sigma}}_{x}\hat{{\bm{B}}}+\frac{1}{l}\left(v_{22}^{2}\sigma^{2}+\operatorname{Tr}(\hat{{\bm{B}}}{\bm{\Sigma}}_{x})\right){\bm{I}}\right){\bm{\Sigma}}_{x},\quad{\bm{F}}_{2}:=({\bm{\mu}}_{w}{\bm{v}}_{21}^{T}+v_{22}{\bm{B}}){\bm{\Sigma}}_{x},$ (13)
@@
The generalization error is derived using Isserlis’ theorem (Isserlis, 1918) to compute higher-order moments. See Appendix G for the full derivation.
@@
Although temperature can be implicitly encoded in $\bm{M}$} during pretraining, it becomes especially important under distribution shifts that the model is not equipped to handle. In such cases, one can optimize generalization performance by choosing the temperature $\tau_{\text{optimal}}$ that minimizes the generalization error, as discussed next.
4.4 Optimal attention temperature improves performance

### v2

# Optimal Attention Temperature Improves the Robustness of In-Context Learning under Distribution Shift in High Dimensions

@@ Abstract
Pretrained Transformers can perform in-context learning (ICL) from a few demonstrations, but this ability can fail sharply when the test distribution differs from pretraining—a common deployment setting. We study attention temperature as a simple inference-time control for improving ICL robustness under such shifts. In a high-dimensional linear-regression framework, we analyze a Transformer with “approximate softmax” attention, which preserves softmax’s normalization and temperature-dependent selectivity while remaining tractable. We derive a closed-form expression for the ICL generalization error under distribution shift, and show that it is minimized by an explicit optimal attention temperature. This characterization yields interpretable guidance by linking the best temperature to moments of the pre-softmax attention scores, and predicts when temperature adjustment can recover near Bayes-optimal performance. We validate the theory with extensive simulations, and further demonstrate gains on pretrained LLMs (GPT-2 and Llama2-7B) on question-answering benchmarks under distribution shift induced by noisy in-context demonstrations. Overall, attention temperature emerges as a principled, lightweight knob for improving the robustness of ICL in pretrained Transformers.
Machine Learning, ICML
@@ 1 Introduction
where $\bm{Z}$} denotes the input representation and $\bm{Q}$}, $\bm{K}$}, and $\bm{V}$} are the query, key, and value weight matrices, respectively. The parameter $\tau>0$, referred to as the attention temperature, controls the variance of the softmax outputs and hence the selectivity of attention weights. This quantity is distinct from the “sampling temperature” commonly used to adjust the output distribution of generative models such as large language models (LLMs) (Renze and Guven, 2024). Throughout this work, we exclusively focus on the attention temperature as an intrinsic component of the attention mechanism.
@@
In this paper, we present a unified theoretical and empirical study of the attention temperature in the context of in-context learning (ICL). We focus on how adjusting this parameter can systematically improve the ICL performance of pretrained Transformers under distributional shift. We address this question in the setting of linear regression tasks, which offer a well-controlled yet expressive framework for dissecting the mechanisms of ICL (Garg et al., 2022; Zhang et al., 2024). Departing from prior work restricted to linear attention, we analyze a Transformer with approximate softmax attention—an architecture that preserves the essential temperature-dependent behavior of standard attention while remaining mathematically tractable.
@@
    We provide, to our knowledge, the first theoretical characterization of the optimal attention temperature for pretrained Transformers with approximate softmax attention in ICL tasks.
@@ 2 Related work
Simplified Transformer variants—particularly those using linear attention—have proven useful for gaining analytical insights about ICL (Garg et al., 2022; Zhang et al., 2024; Raventós et al., 2023). Notably, Zhang et al. (2024) showed that linear Transformers approximate Bayes-optimal inference in linear regression tasks, even under distribution shift. We build on this line of research but focus explicitly on the role of the attention temperature. In contrast to Zhang et al. (2024), we (i) employ approximate softmax attention to isolate the effect of temperature, (ii) study how temperature adjustments can mitigate the impact of distribution shifts, and (iii) derive and empirically evaluate the optimal temperature for improving ICL performance. These advances extend prior analyses and yield a deeper theoretical and empirical understanding of how principled temperature selection enhances the robustness of Transformers under distributional shift. [^1] Due to space limitations, additional related work is discussed in Appendix A.
@@
Although linear attention has gained traction for its computational efficiency, it typically lags behind softmax-based counterparts in predictive performance, spurring efforts to narrow this gap (Choromanski et al., 2021; Qin et al., 2022). A pivotal advance in this direction is due to Han et al. (2024), who showed that a approximate variant of softmax attention can closely approximate the performance of standard softmax attention. Building on this insight, we adopt the approximate softmax formulation, which preserves the essential temperature-dependent behavior of standard attention while enabling tractable theoretical analysis. This choice provides a principled framework for investigating how attention-temperature selection shapes ICL performance in pretrained Transformers.
@@
Figure 1: Comparison of temperature effects in softmax, approximate softmax, and linear (with temperature scaling) cases. We consider an input vector ${\bm{x}}\in\mathbb{R}^{l}$ whose histogram is illustrated on the left-most plot. Rest of the plots illustrates histograms of the elements of $\text{softmax}({\bm{x}}/\tau)$, approximation $\widehat{\text{softmax}}({\bm{x}}/\tau)$ derived in Appendix C and ${\bm{x}}/(l\tau)$ from left to right, respectively.
@@ 3 Setting
We describe the setup for analyzing ICL in linear regression using pretrained Transformers, covering the data model, approximate softmax attention with reparameterization, evaluation metrics, and the Bayes-optimal benchmark.
3.1 In-context learning for linear regression
@@
${\bm{x}}_{i}\sim{\mathcal{N}}({\bm{\mu}}_{x},{\bm{\Sigma}}_{x}),\;y_{i}={\bm{w}}^{T}{\bm{x}}_{i}+\epsilon_{i},\;\epsilon_{i}\sim{\mathcal{N}}(0,\sigma^{2}),$ (2)
@@
$\|{\bm{\mu}}_{x}\|,\|{\bm{\mu}}_{w}\|\leq c_{1},\quad\lambda_{\min}({\bm{\Sigma}}_{x}),\lambda_{\min}({\bm{\Sigma}}_{w})\geq c_{2},$ (3)
$\lambda_{\max}({\bm{\Sigma}}_{x}),\lambda_{\max}({\bm{\Sigma}}_{w})\leq c_{3}.$ (4)
@@ 3.2 Modeling attention with transformers
3.3 Approximate softmax attention
To analytically characterize the effect of temperature on ICL, we employ an approximation of softmax attention :
${\bm{E}}:={\bm{Z}}+{\bm{V}}{\bm{Z}}\cdot\widehat{\text{softmax}}\left(\frac{({\bm{K}}{\bm{Z}})^{T}({\bm{Q}}{\bm{Z}})}{\tau}\right),$ (7)
the explicit form of which is provided together with its derivation in Appendix C. In contrast to linear attention (Zhang et al., 2024),
@@ 3.3 Approximate softmax attention
Remark 3.4 (Linear vs. approximate softmax attention).
Approximate softmax attention maintains row-wise normalization, making it inherently more robust to shifts in input means — a critical failure mode of linear attention in ICL.
Another key distinction between the linear case and the approximate softmax case is that linear (with temperature scaling) fails to capture the temperature behavior of softmax. However, while this may not be immediately apparent, approximate softmax closely mirrors the behavior of softmax with respect to temperature variation. Figure 1 visualizes the effects of temperature across softmax, approximate softmax, and linear (with temperature scaling) settings under an example input. The figure demonstrates our claim that approximate softmax can be a useful tool to study the temperature effect exhibited by softmax. We expand on this point in Appendix D.
3.4 Reparametrization of approximate softmax attention
@@
By analyzing this reparameterization, we gain a deeper understanding of how the model parameters interact with the data to address the ICL problem effectively.
@@ 3.6 Bayes-optimal ridge estimator
The derivation, provided in Appendix B, illustrates how Bayesian principles inform regression by integrating data evidence with prior distributions to yield more reliable predictions. In our setting, the inputs and labels are derived from the prompt matrix $\bm{Z}$}, and the Bayes-optimal linear model predicts any input $\bm{x}$} as $\hat{{\bm{w}}}_{Bayes}^{T}{\bm{x}}$.
@@
4 Theoretical results
In this section, we present our main theoretical results on the ICL under distribution shifts for the Transformer with an approximate softmax attention without MLP layers, denoted by (7). We provide an asymptotic characterization of its generalization error, offering a principled framework for analyzing performance. Next, we investigate the role of the temperature parameter and demonstrate that adjusting it appropriately can substantially improve generalization—especially in cases where the model initially fails to perform effective in-context learning. We then identify specific conditions for which the model fails to generalize under distribution shifts at test time, revealing key limitations of the model in ICL. We describe the interpretation of the found analytical form of the optimal attention temperature in mitigating the adverse effects of distribution shift. Finally, we explain a possible extension of our optimal temperature to other ICL settings involving softmax attention.
@@ 4.1 In-context learning performance
We analyze the in-context learning (ICL) performance of the Transformer model (7) by evaluating the generalization error defined in (10). To establish a general setting for the subsequent results, we impose the following assumption on the pretrained parameters.
@@
The generalization error result stated below holds for any parameters ${\bm{M}},{\bm{V}}$ that satisfy Assumption 4.1.
@@
Suppose Assumptions 3.13.2 and 4.1 hold. At test time, assume the input, task, and noise distributions are given by $\mathcal{N}({\bm{\mu}}_{x},{\bm{\Sigma}}_{x})$, $\mathcal{N}({\bm{\mu}}_{w},{\bm{\Sigma}}_{w})$, and $\mathcal{N}(0,\sigma^{2})$, respectively. Define
@@
$\mathcal{G}({\bm{V}},{\bm{M}})$ $=\frac{1}{\tau^{2}}\operatorname{Tr}\left({\bm{A}}{\bm{M}}_{11}^{T}{\bm{F}}_{1}{\bm{M}}_{11}\right)$
$\quad-\frac{1}{\tau}\operatorname{Tr}\left({\bm{A}}\left({\bm{F}}_{2}{\bm{M}}_{11}+{\bm{M}}_{11}^{T}{\bm{F}}_{2}^{T}\right)\right)$
$\quad+\operatorname{Tr}\left({\bm{A}}{\bm{B}}\right)+\sigma^{2},$ (12)
@@
${\bm{F}}_{1}$ $:=\left({\bm{\Sigma}}_{x}\hat{{\bm{B}}}+\frac{1}{l}\left(v_{22}^{2}\sigma^{2}+\operatorname{Tr}(\hat{{\bm{B}}}{\bm{\Sigma}}_{x})\right){\bm{I}}\right){\bm{\Sigma}}_{x},$ (13)
${\bm{F}}_{2}$ $:=({\bm{\mu}}_{w}{\bm{v}}_{21}^{T}+v_{22}{\bm{B}}){\bm{\Sigma}}_{x},$ (14)
@@
The generalization error is derived using Isserlis’ theorem (Isserlis, 1918) to compute higher-order moments. See Appendix G for the full derivation. ∎
@@
Although temperature can be implicitly encoded in $\bm{M}$} during pretraining, it becomes especially important under distribution shifts that the model is not equipped to handle. In such cases, one can optimize generalization performance by choosing the temperature $\tau_{\text{opt}}$ that minimizes the generalization error, as discussed next.
4.2 Optimal attention temperature